# Without Paired Labeled Data: An End-to-End Self-Supervised Paradigm for UAV-View Geo-Localization

### v2

@@ III-A Dual-Path Contrastive Learning Baseline
Cluster-based contrastive learning methods have achieved remarkable success in various fields. However, in UVGL, due to the unique challenges of cross-view scenarios, single-path cluster-based contrastive learning struggles to effectively capture view-specific features. To address this issue, we design the baseline network with a dual-path architecture, where UAV-view and satellite-view features are extracted using a shared backbone and clustered to assign pseudo-labels. Subsequently, a dual-path contrastive learning strategy [53] is employed to enhance intra-view representation learning.
We first introduce the notations used in this paper for better clarity. Let denote a set of drone images with instances, and denote a set of satellite images with instances. and represent the feature representations extracted by the shared feature extractor for the drone and satellite views, respectively. and denote the drone query instance features and satellite query instance features extracted by the feature extractor . The processing procedure of the backbone can be described as
Cross-View Feature Memory Initialization. At the beginning of each training epoch, the feature memories for both views are initialized. Specifically, the cluster representations of the drone view and the satellite view are stored in the drone memory dictionary and satellite memory dictionary , respectively, for view-specific memory initialization. This process can be formalized as
where , , , , denotes the -th cluster set in drone or satellite view, indicates the number of instances per cluster.
Cross-View Feature Memory Updating. During the training phase, we randomly sample distinct localization points from the training set, with each point comprising a fixed number of instances. This sampling strategy results in a minibatch containing a total of query images. Subsequently, the query features extracted from the minibatch are integrated into the cluster representations via a momentum-based update mechanism. Through this method, we iteratively update the features for both perspectives. This process can be formalized as
where is the momentum updating factor. is the iteration number.
Cross-View Contrastive Loss. Given drone and satellite query and , we compute the contrastive loss for drone view and satellite view by the following equations:
where denotes the temperature hyper-parameter. The cluster centroids are utilized as feature vectors at the cluster level to compute the distances between the query instance and all clusters. Here, corresponds to the positive cluster feature associated with .
The final optimization for baseline is denoted by the following combination:
@@ III-B Dynamic Hierarchical Memory Learning
Geographic environments inherently exhibit spatial continuity, where neighboring regions usually present similar visual patterns, while distant regions display significant differences. However, such spatial characteristics often introduce ambiguity in intra-view feature representations. To address this, we propose a DHML module that jointly exploits long-term and short-term memory to capture spatial dynamics and enhance the discriminability of intra-view features.
Consistent with the baseline network, we construct the dynamic hierarchical memory learning in a dual-path manner, using drone feature extraction as an illustrative example. To build the long-term memory, we first construct two novel clustering memory dictionaries according to Eq.2, which are used to store the short-term and long-term memories, respectively.
where, represent the short-term memory embeddings that encode the cluster representations of the drone view, and represent the long-term memory embeddings that store historically accumulated stable representations of the drone view.
Unlike the weight updating scheme in Eq.4 and Eq.5, we assign different weights to the input features and the historical memory during the long-term memory update. Specifically, since the momentum parameter is set to 0.2 throughout this study, we enhance the influence of historical memory during the initialization of long-term memory updates. The update procedure can defined as
where, represents the drone query instance features. This obtain a new dictionary of long-term memory .
Additional, we further introduce an adaptive weight updating mechanism to update the long-term memory. In detail, we compute the Euclidean distance between the current input and its corresponding long-term memory, then take the average over all samples in the batch. This average distance is then mapped to the range via the sigmoid function, yielding an adaptive update coefficient :
where, denotes the batch size and represents the sigmoid function. The short-term memory is subsequently updated as follows:
@@
where denotes the temperature hyper-parameter. The combined cluster centroids are utilized as feature vectors at the cluster level to compute the distances between the query instance and all clusters. Here, corresponds to the positive cluster feature associated with .
Finally, we introduce a fusion factor to perform weighted fusion of and , aiming to enhance the discriminability of intra-view feature representations and obtain the optimized loss . Similarly, can be obtained in the same way. Therefore, the dynamic hierarchical memory learning process can be summarized as
@@ III-C Information Consistency Evolution Learning
The baseline network and the DHML module primarily focus on intra-view feature learning, lacking the ability to model and associate cross-view features effectively. Therefore, the proposed ICEL module employs a threshold-based and ranking-based neighborhood selection strategy to achieve effective cross-view feature alignment. By promoting the consistency evolution of cross-view feature representations, ICEL enhances both the discriminability and representation capability of cross-view features.
@@
where and denote the drone and satellite instance features stored in memory dictionaries and , respectively.
In addition, we define the similarity between a given drone (or satellite) query and each drone (or satellite) instance in the training set as
where and are the indices of queries and instances. Thus, we obtain intra-view similarity , and cross-view similarity , .
@@
where represents the set of drone feature instances selected as valid intra-view neighbors. denotes the set of drone feature instances that exhibit strong cross-view associations with satellite features. is a predefined threshold. The and can be obtained by the same way.
@@
where , and , denote the selected intra-view and cross-view neighborhood sets, respectively, with and containing exactly elements, and and containing exactly elements. Similarly, the neighborhood sets , , , and can be obtained following the same selection strategy.
@@
where represents the batch size of the query , and denote the total number of neighborhood samples within the set and .
@@
where is the softmax-normalized similarity distribution, is used for summation to normalize the similarity scores over all neighborhood samples. and can be obtained in the same way.
Within a more restrictive and smaller set of neighbors, a mutual information constraint is proposed to learn intra-view and cross-view features representations. This complements the consistency constraint described above, achieving a balance between global structural alignment and local discriminative information. Within the selected neighborhood, we first compute, for each neighbor sample (where ), the similarity between the deep query feature and the neighbor feature using the similarity function . Based on these similarity scores, the posterior probability for each neighbor is defined via a softmax normalization as:
where denote the total number of neighborhood samples within the set and .
Therefore, the intra-view and cross-view mutual information losses and can be derived and are defined as follows:
where and represent the normalized similarity distributions over the selected intra-view and cross-view neighbors, respectively. and can be obtain in the same way.
For each view combination, the loss is computed as the weighted sum of three components: the original neighborhood loss , the strict neighborhood loss , and the extended neighborhood loss :
where and denote the drone or satellite view.
The total neighbor learning loss is then formulated as
@@ III-D Pseudo-Label Enhancement
In self-supervised learning, the quality of pseudo-labels exerts a decisive influence on model performance. Inspired by [52], our proposes a pseudo-label optimization strategy based on dynamic feature robustness. Specifically, we introduce small perturbations to the original feature representations and exploit the top-K neighbor consistency between the original and perturbed features to identify high-confidence samples while filtering out noisy instances. Then, a spatial smoothing operation based on the similarity matrix is applied to refine the pseudo-labels, enforcing local label consistency.
We first employe small Gaussian noise into the original satellite instances feature and the original drone instances features to obtain perturbed features, which are used for more robust similarity estimation and subsequent cluster consistency evaluation. It can be formulated as
where, with the similarities and , we can get two cross-view ranking lists and for and , which are denoted as and .
The label of the -th similar infrared instance in two ranking lists can be represented as
where and are refined -th cross-view labels of and through the ranking on cross-view initial features and perturbed features similarities.
So, we associate the cross-modality labels with the intersection of two label sets to investigate collaborative ranking consistency. This process can be described as
where records the samples with the same identity of top- identity in instances from top-1 to top-.
The reliable refined cross-view label of can be expressed by the label of maximum count in .
where denotes the counting function. The operation traverses and finds the labels with maximum number as cross-view refined labels.
Then, we convert the label list to the form of one-hot code matrix by setting the column according to the refined labels to 1 and the rest to 0.Then,two initial and perturbed homogeneous initial and deep similarity matrices are constructed to investigate intra-view ranking consistency, enhancing the precision of refined cross-view pseudo labels by:
where represents the intra-view similarity structure.
In order to explore the intra-view initial and perturbed collaboration, we calculate the sum of and by:
where indicates the consistency of initial and perturbed similarity matrix.
We keep the 5-max values of in each row to 1 and set the rest to 0, acquiring the ranking relations. The process of intra-modality ranking smoothing is formulated as follows:
where is the final refined cross-view label matrix of satellite instance. In , the column number of the maximum value in each row is the refined label of samples.
@@
| Model | Venue | Supervision Ratio | Paramars(M) | R@1 | AP | R@1 | AP |
@@ IV-A Experimental Datasets and Evaluation Metrics
To evaluate our self-supervised method, we conduct experiments on three benchmark datasets including University-1652 [12], SUES-200 [33] and DenseUAV [54].
University-1652 contains images from 1,652 buildings across 72 universities worldwide. It includes drone, satellite, and ground-level views. The training set consists of 701 buildings from 33 universities, and the test set covers 951 buildings from 39 geographically disjoint universities.
SUES-200 focuses on altitude diversity in aerial imagery, containing 200 locations with 120 for training and 80 for testing. Each location provides one satellite image and aerial images captured at four altitudes (150m, 200m, 250m, and 300m), covering urban areas, natural landscapes, and water bodies. SUES-200 evaluates cross-view retrieval under varying flight heights, reflecting real-world UAV scenarios.
DenseUAV focuses on UAV self-positioning in low-altitude urban environments, containing over 27,000 UAV and satellite images collected from 14 university campuses. It features dense sampling, multi-scale satellite images, and varying altitudes (80m, 90m, and 100m) under multiple temporal conditions. The training set includes 6,768 UAV images and 13,536 satellite images, while the test set comprises 2,331 UAV images and 4,662 satellite images from non-overlapping areas. DenseUAV poses challenges such as cross-view matching, neighbor confusion, and spatial-temporal variation.
We evaluate the models using Recall@K (R@K) and Average Precision (AP), where R@K measures the proportion of correct matches among the Top-results and AP balances precision and recall. Model efficiency is assessed by the number of parameters, reflecting portability under resource constraints. All comparisons are conducted using the best configurations for each method.
@@ IV-B Implementation Details
A ConvNeXt-Tiny model, pre-trained on ImageNet, serves as the backbone for feature extraction, with a newly added classifier module initialized via the Kaiming method. During both training and testing, all input images are resized to . We employ several data augmentation techniques, including random cropping, random horizontal flipping, and random rotation. The batch size is set to 64, and DBSCAN [8] is used to generate pseudo labels. For optimization, we use the SGD optimizer with an initial learning rate of 0.001, training the model for 30 epochs in total. All experiments are conducted under the PyTorch framework on an Ubuntu 22.04 system equipped with four NVIDIA RTX 4090 GPUs. Additionally, in the University-1652 and SUES-200 datasets,we replicate each satellite feature 50 times for clustering purposes, leaving UAV features unchanged. In DenseUAV, we replicate UAV features four times at each altitude, and replicate the satellite features three times for each time point and altitude.
@@ IV-C Comparison with State-of-the-art Methods
To evaluate the effectiveness of our method, we compare it with the state-of-the-art supervised methods MuSe-Net [54], MCCG [1], SDPL [3], CCR [32], Sample4Geo [29], SRLN[31], MEAN [2] and DAC [4], and the recent self-supervised methods EM-CVGL [5], CDIKTNet [42].
Results on University-1652. As shown in Table I, our method exhibits superior performance compared with the current advanced self-supervised methods. Specifically, our method achieves R@1/AP scores of 90.17%/91.67% and 95.01%/88.95% under the DroneSatellite and SatelliteDrone settings, respectively. Compared with EM-CVGL, our method achieves significant improvements, outperforming it by approximately 20% in R@1 and 17% in AP on the DroneSatellite setting, and by 16% in R@1 and 28% in AP on the SatelliteDrone setting. Furthermore, even compared with CDIKTNet, which utilizes paired data from other domains for self-supervised fine-tuning, our method still consistently achieves superior performance without relying on any cross-domain paired data.
We further benchmark against eight representative supervised methods, and the results show that our method clearly surpasses some of them, including MuSe-Net, MCCG, and SDPL. Furthermore, with self-supervised pre-training and subsequent supervised fine-tuning, our method achieves competitive performance under the supervised setting, comparable to CCR, Sample4Geo, and SRLN.
@@
| Model | Venue | Supervision Ratio | Paramars(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
@@
| Model | Venue | Supervision Ratio | Paramars(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
@@
Results on SUES-200. As shown in Table II, As shown in Table II, our method, when directly transferring the pre-trained model from the University-1652 dataset to the SUES-200 dataset without any re-training, consistently outperforms all existing self-supervised methods and achieves competitive performance compared to supervised methods such as MCCG and SDPL. Moreover, by performing self-supervised re-training on the SUES-200 dataset, our method brings a significant performance boost, achieving results comparable to the supervised CCR. Notably, when further fine-tuning the transferred model for only one epoch, we observe additional improvements, surpassing the supervised SRLN. These results clearly demonstrate the effectiveness of our method in self-supervised learning, transfer learning, and efficient fine-tuning.
Results on Cross-Domain Generalization Performance: Our method consistently outperforms existing self-supervised method under cross-domain settings, showing both significant improvements and stable generalization. In the Drone to Satellite setting, our model achieves better performance than EM-CVGL at all altitudes. At 150 meters, R@1 and AP increase by approximately 22% and 20%; at 200 meters, the gains are around 19% and 16%; at 250 meters, nearly 14% and 12%; and even at 300 meters, our method still outperforms by about 11% and 9%. In the Satellite to Drone scenario, similar improvements are observed. At 150 meters, R@1 and AP improve by approximately 18% and 25%; at 200 meters, by 10% and 20%; at 250 meters, by 6% and 13%; and at 300 meters, by 3% and 10%. These results confirm the robustness and strong cross-domain generalization capability of our method across varying altitudes.Remarkably, our method even surpasses all existing state-of-the-art supervised methods in terms of cross-domain generalization. These results confirm the robustness and strong cross-domain generalization capability of our method across varying altitudes.
@@
| Model | Venue | Supervision Ratio | Paramars(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
@@
| Model | Venue | Supervision Ratio | Paramars(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
@@
| Model | Venue | Supervision Ratio | Paramars(M) | R@1 | R@5 | R@1 | R@5 | R@1 | R@5 | R@1 | R@5 |
@@
Results on DenseUAV: The DenseUAV dataset presents a highly challenging scenario due to its densely sampled nature, where neighboring images exhibit high visual similarity, often leading to severe confusion in UVGL. In addition, the dataset includes satellite images captured at different times and UAV images from multiple altitudes, which further increases intra-domain variability and task complexity. As shown in Table IV, under the fully unsupervised setting, our method achieves 73.79% R@1 and 90.69% R@5 across all altitudes, significantly outperforming existing self-supervised methods. It is noteworthy that the two representative self-supervised baselines consistently yield poor performance on DenseUAV, highlighting their limitations in handling densely sampled and spatiotemporal diverse scenarios. Furthermore, after supervised fine-tuning, our method surpasses several supervised baselines, including ConvNeXt-Tiny, Sample4Geo, CAMP, and DAC. In some cases, it even outperforms the current best-performing supervised method, MEAN. These results clearly demonstrate the effectiveness of our method in addressing extreme feature confusion caused by spatial continuity and appearance shifts induced by temporal variation, confirming its strong generalization and robustness.

### v3

@@ III-A Dual-Path Contrastive Learning Baseline
Cluster-based contrastive learning methods have achieved remarkable success in various fields [17, 50]. However, in DVGL, the unique challenges of cross-view scenarios, as discussed in Section I, make it difficult for single-path cluster-based contrastive learning to effectively capture view-specific features. Therefore, we design the baseline network with a dual-path architecture, where drone-view and satellite-view features are extracted using a shared backbone and clustered to assign pseudo-labels. Subsequently, a dual-path contrastive learning strategy [51] is employed to enhance intra-view representation learning.
We first introduce the notation for better clarity. Let denote a set of drone images with instances, and denote a set of satellite images with instances. and represent the feature representations extracted by the shared feature extractor for drone-view and satellite-view, respectively. and denote the drone query instance features and satellite query instance features extracted by the feature extractor .
Arbitrary Sampling and Backbone-based Feature Extraction. To simulate real-world scenarios where drone and satellite images are not paired, we independently and arbitrarily sample drone and satellite images at each training epoch to construct unpaired batches. Each image is subsequently processed by a shared backbone network to extract features. The backbone processing procedure can be described as
Cross-View Cluster Memory Initialization. At the beginning of each training epoch, the feature memories for both views are initialized into clusters and clusters, respectively. Specifically, the cluster centroids of the drone-view and the satellite-view are stored in the drone memory dictionary and the satellite memory dictionary , respectively, for view-specific memory initialization. This process can be formalized as
where denotes the -th cluster set in drone-view, , and denotes the -th cluster set in satellite-view, . indicates the number of instances per cluster.
Cross-View Cluster Memory Updating. During the training phase, we randomly sample different localization points from the training set, with each point comprising a fixed number of instances. This sampling strategy results in a minibatch containing a total of query images. Subsequently, the query features extracted from the minibatch are integrated into the cluster representations via a momentum-based update mechanism. The cluster centroids and are updated iteratively as
where is the momentum updating factor.
Cross-View Cluster Memory Contrastive Loss. Given drone and satellite query and , we compute the contrastive loss for drone-view and satellite-view as
where denotes the temperature hyper-parameter. Cluster centroids and are utilized as feature vectors at the cluster level to compute the distances between the query instance and all clusters. Here, and are positive cluster features associated with and . The final optimization for baseline is denoted as
@@ III-B Dynamic Hierarchical Memory Learning
Geographic environments inherently exhibit spatial continuity, where neighboring regions usually present similar visual patterns, while distant regions show significant differences. However, such spatial characteristics often introduce ambiguity in intra-view feature representations. To address this issue, we propose a DHML module that jointly exploits long-term and short-term memory to capture spatial dynamics and enhance the discriminability of intra-view features.
Consistent with the baseline network, we construct dynamic hierarchical memory learning in a dual-path manner, using drone feature extraction as an illustrative example. To build long-term memory, we first construct two novel clustering memory dictionaries according to Eq.(2), which are used to store short-term and long-term memories, respectively. They are expressed as
where represent the short-term memory embeddings that encode the cluster representations of the drone-view, and represent the long-term memory embeddings that store historically accumulated stable representations of the drone-view.
Unlike the weight updating scheme in Eq.(4), we assign different weights to the input features and historical memory during the long-term memory update. Specifically, since the momentum parameter is set to 0.2 throughout this study, we enhance the influence of historical memory during the initialization of long-term memory updates. The update procedure can defined as
where represents the drone query instance features. This obtains a new dictionary of long-term memory.
In addition, we further introduce an adaptive weight update mechanism to update the long-term memory. In detail, we compute the euclidean distance between the current input and its corresponding long-term memory, then take the average over all samples in the batch. This average distance is then mapped to the range via the sigmoid function, yielding an adaptive update coefficient:
where denotes the batch size and represents the sigmoid function. The short-term memory is subsequently updated as
@@
where denotes the temperature hyper-parameter. The combined cluster centroids are used as feature vectors at the cluster level to calculate the distances between the query instance and all clusters. Here, corresponds to the positive cluster feature associated with . Finally, we introduce a fusion factor to perform a weighted fusion of and , aiming to enhance the discriminability of intra-view feature representations and obtain the optimized loss . Similarly, can be obtained in the same way. Therefore, the dynamic hierarchical memory learning process can be summarized as
@@ III-C Information Consistency Evolution Learning
The baseline network and the DHML module primarily focus on intra-view feature learning, lacking the ability to model and associate cross-view features effectively. Therefore, the proposed ICEL module employs a threshold-based and ranking-based neighborhood selection strategy to achieve effective cross-view feature alignment. By promoting the consistency evolution of cross-view feature representations, ICEL enhances both the discriminability and the representation capability of cross-view features.
@@
where , , and , , denote the drone and satellite instance features stored in memory dictionaries and , respectively.
To facilitate the following description, we take the scenario where the drone-view serves as the query as an illustrative example. The similarity between a given drone query that denotes the -th query sample in the drone-view, . denote the batch size of the query . Each drone instance and each satellite instance in the training set are compared to drone query by computing a similarity score, respectively. This is defined as
@@
where represents the set of drone feature instances selected as valid intra-view neighbor, and denotes the set of drone feature instances that exhibit strong cross-view associations with satellite features. is a predefined threshold.
@@
where and denote the selected intra-view and cross-view neighborhood sets, respectively. and can be obtained following the same selection strategy.
@@
where and denote the total number of neighborhood samples within the set and .
@@
where and are the softmax-normalized similarity distribution.
Within a more restrictive and smaller set of neighbors, a mutual information constraint is proposed to learn intra-view and cross-view features representations. This complements the consistency constraint described above and achieves a balance between global structural alignment and local discriminative information. Within the selected neighbors, we compute the similarity between the query and each neighbor and using the function , and apply softmax to obtain the posterior probability for each neighbor. It can be described as
Therefore, the intra-view and cross-view mutual information losses and can be derived and are defined as
For each view, the loss is formulated as a weighted sum of the original neighborhood loss , the strict neighborhood loss , and the extended neighborhood loss . Specifically, the intra-view loss and the cross-view loss are respectively defined as
where and are hyperparameters controlling the contribution of strict and extended neighborhood components. When the satellate view as the query, the and can be obtained in the same way. The total neighbor learning loss is computed as .
@@ III-D Pseudo-Label Enhancement
In self-supervised learning, the quality of pseudo-labels exerts a decisive influence on model performance. Inspired by [49], we propose a pseudo-label optimization strategy based on dynamic feature robustness. Specifically, we introduce small perturbations to the original feature representations and exploit the top-K neighbor consistency between the original and perturbed features to identify high-confidence samples while filtering out noisy instances. Then, a spatial smoothing operation based on the similarity matrix is applied to refine the pseudo-labels to enhance local label consistency.
Small Gaussian noise is used in the initial satellite instances feature and the initial drone instances features to obtain perturbed features, where the denotes the standard deviation. It can be formulated as
Using the definition of the similarity score shown in Eq.(16), two cross-view ranking lists can be obtained, which are denoted as and , . The label of the -th similar drone instance in two ranking lists can be denoted as and . Next, the cross-view labels are associated with the intersection of two label sets to investigate collaborative ranking consistency. This process can be described as
where records the samples with the same identity in instances. Accordingly, we update the cross-view pseudo-label for using the label with the highest frequency in , which reflects the most stable identity under both original and perturbed feature rankings. This final refined label is denoted as . Then, we convert the label list into the form of one-hot code matrix by setting the column according to the refined labels to 1 and the rest to 0.
In order to explore the intra-view initial and perturbed collaboration, the initial and perturbed similarity matrices and are calculated, and their sum is used to investigate intra-view ranking consistency. Then we keep the 5-max values of in each row at 1 and set the rest to 0, acquiring the ranking relations. The process of intra-view ranking smoothing is formulated as
where is the final refined cross-view label matrix of satellite instance. In , the column number of the maximum value in each row is the refined label of the samples.
@@
Model | Venue | Supervision Ratio | Parameters(M) | R@1 | AP | R@1 | AP |
@@ IV-A Experimental Datasets and Evaluation Metrics
To evaluate the proposed method, we conduct experiments on three benchmark datasets, including University-1652 [12], SUES-200 [32] and DenseUAV [52].
University-1652 contains abundant images of 1,652 buildings in 72 universities around the world. It involves drone, satellite, and ground-level views. The training set consists of 701 buildings from 33 universities and the test set covers 951 buildings from 39 geographically nonaffiliated universities.
SUES-200 focuses on altitude diversity in aerial imagery, containing 200 locations, with 120 designated for training and 80 for testing. Each location in the dataset contributes one satellite image along with aerial images captured at four distinct altitudes, which involve urban areas, natural landscapes, and water bodies. SUES-200 evaluates cross-view retrieval under varying flight heights, reflecting real-world drone scenarios.
DenseUAV focuses on drone self-positioning in low-altitude urban environments with more than 27,000 drone and satellite images collected from 14 university campuses. It is characterized by dense sampling, multi-scale satellite imagery, and diverse altitudes captured across multiple temporal conditions. DenseUAV poses challenges such as cross-view matching, neighbor confusion, and spatial-temporal variation.
We evaluate the models using Recall@K (R@K) and Average Precision (AP), where R@K measures the proportion of correct matches among the Top-results and AP evaluates the overall precision-recall performance. Model efficiency is assessed by the number of parameters to reflect portability under resource constraints. All comparisons are conducted using the best configurations for each method.
@@ IV-B Implementation Details
A ConvNeXt-Tiny model, pre-trained on ImageNet, serves as the backbone for feature extraction, with a newly added classifier module initialized via the Kaiming method. During both training and testing, all input images are resized to . We employ several data-augmentation techniques, including random cropping, random horizontal flipping, and random rotation. The batch size is set to 64, and DBSCAN [7] is used to generate pseudo-labels. For optimization, we use the SGD optimizer with an initial learning rate of 0.001, training the model for 30 epochs in total. All experiments are conducted under the PyTorch framework on an Ubuntu 22.04 system equipped with four NVIDIA RTX 4090 GPUs. Additionally, for the University-1652 and SUES-200 datasets, we apply feature replication only on the training sets by duplicating each satellite feature 50 times for clustering purposes, while drone features remain unchanged. The test sets are kept intact without any modification. In the DenseUAV dataset, drone features are replicated four times at each altitude, and satellite features are replicated three times for each time point and altitude exclusively in the training set; the test set remains unchanged.
@@ IV-C Comparison with State-of-the-art Methods
To evaluate the effectiveness of our method, we compare it with the state-of-the-art supervised methods MuSe-Net [52], MCCG [1], SDPL [53], CCR [31], Sample4Geo [28], SRLN [30], MEAN [2] and DAC [3], and the self-supervised methods EM-CVGL [4], CDIKTNet [40] and Wang [9].
Results on University-1652. As shown in Table I, our method exhibits a superior performance compared to current advanced self-supervised methods. Specifically, our method achieves R@1/AP scores of 90.17%/91.67% and 95.01%/88.95% under the DroneSatellite and SatelliteDrone settings, respectively. Compared to EM-CVGL, our method achieves significant improvements, outperforming it by approximately 20% in R@1 and 17% in AP on the DroneSatellite setting, and by 16% in R@1 and 28% in AP on the SatelliteDrone setting. Furthermore, even compared to CDIKTNet, which utilizes paired data from other domains for self-supervised fine-tuning, our method still consistently achieves superior performance without relying on any cross-domain paired data. Furthermore, compared with the latest clustering-based self-supervised methods, our approach still achieves significant improvements.
We further benchmark with eight representative supervised methods, and the results show that our method clearly surpasses some of them, including MuSe-Net, MCCG, and SDPL. Furthermore, with self-supervised pre-training and subsequent supervised fine-tuning, our method achieves competitive performance under the supervised setting, comparable to CCR, Sample4Geo, and SRLN.
@@
Model | Venue | Supervision Ratio | Parameters(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
@@
Model | Venue | Supervision Ratio | Parameters(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
@@
Results on SUES-200. As shown in Table II, our method, when directly transferring the pre-trained model from the University-1652 dataset to the SUES-200 dataset without any re-training, consistently outperforms all existing self-supervised methods and achieves competitive performance compared to supervised methods such as MCCG and SDPL. Moreover, by performing self-supervised re-training on the SUES-200 dataset, our method brings a significant performance boost, achieving results comparable to the supervised CCR. Notably, when further fine-tuning the transferred model for only one epoch, we observe additional improvements, surpassing the supervised SRLN. These results clearly demonstrate the effectiveness of our method in self-supervised learning, transfer learning, and efficient fine-tuning.
Results on Cross-Domain Generalization Performance. As shown in Table III, our method consistently outperforms existing self-supervised methods in cross-domain settings, showing both significant improvements and stable generalization. Specifically, in the Drone to Satellite setting, our model achieves better performance than EM-CVGL at all altitudes. At 150 meters, R@1 and AP increase by approximately 22% and 20%. At 200 meters, the gains are around 19% and 16%; at 250 meters, nearly 14% and 12%. Even at 300 meters, our method still outperforms by about 11% and 9%. In the Satellite to Drone scenario, similar improvements are observed. At 150 meters, R@1 and AP improve by approximately 18% and 25%; at 200 meters, by 10% and 20%; at 250 meters, by 6% and 13%; and at 300 meters, by 3% and 10%. These results confirm the robustness and strong cross-domain generalization capability of our method across varying altitudes. Remarkably, our method even surpasses all existing state-of-the-art supervised methods in terms of cross-domain generalization. These results confirm the robustness and strong cross-domain generalization capability of our method across varying altitudes.
@@
Model | Venue | Supervision Ratio | Parameters(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
@@
Model | Venue | Supervision Ratio | Parameters(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
@@
Model | Venue | Supervision Ratio | Parameters(M) | R@1 | R@5 | R@1 | R@5 | R@1 | R@5 | R@1 | R@5 |
@@
Results on DenseUAV. The DenseUAV dataset presents a highly challenging scenario due to its densely sampled nature, where neighboring images exhibit high visual similarity, often leading to severe confusion in DVGL. In addition, the dataset includes satellite images captured at different times and drone images from multiple altitudes, which further increases intra-domain variability and task complexity. As shown in Table IV, in the fully unsupervised setting, our method achieves 73.79% R@1 and 90.69% R@5 across all altitudes, significantly outperforming existing self-supervised methods. It is noteworthy that the two representative self-supervised baselines consistently yield poor performance on DenseUAV, due to their limitations in handling densely sampled and spatio-temporal diverse scenarios. Furthermore, after supervised fine-tuning, our method surpasses several supervised baselines, including ConvNeXt-Tiny, Sample4Geo, CAMP, and DAC. In some cases, it even outperforms the current best-performing supervised method, MEAN. These results clearly demonstrate the effectiveness of our method in addressing extreme feature confusion caused by spatial continuity and appearance shifts induced by temporal variation, confirming its strong generalization and robustness.